OpenStudy (anonymous):

Help, Evaluating Intergrals using substitution prior to intergration by parts.

OpenStudy (anonymous):

\[\int\limits e^\sqrt(3s+9)\]

OpenStudy (bahrom7893):

let a = 3s + 9, then da = 3ds then multiply and divide by 3, to get the 3*ds in da (1/3) Int(e^(sqrt(a)) da)

OpenStudy (anonymous):

but that's not going to help!

OpenStudy (bahrom7893):

yes it will

OpenStudy (bahrom7893):

\[\int\limits_{}^{}e^\sqrt{a}da\]

OpenStudy (bahrom7893):

(1/3) * \[\int\limits_{}^{}e^\sqrt{a}da\]

OpenStudy (anonymous):

what about the u & dv?

OpenStudy (bahrom7893):

now let b^2 = a

OpenStudy (bahrom7893):

da = 2b*db

OpenStudy (bahrom7893):

wait i got lost myself..

OpenStudy (bahrom7893):

hey there sorry i gtg now, but i found how to integrate e^sqrt(x) here: http://answers.yahoo.com/question/index?qid=20080903004829AAGDQnl GOOGLE WORKS WONDERS!!!